OpenStudy (anonymous):

x+5/3≤ 2 solve an show work

OpenStudy (ash2326):

x+5/3<=2 x<=2-5/3 x<=6/3-5/3 x<=1/3

OpenStudy (anonymous):

-5s-8/4\[\ge\]-22

OpenStudy (anonymous):

solve and show work plz

OpenStudy (ash2326):

-5s-8/4>=-22 -5s-2>=-22 -5s>=-20 divide by -5 both sides, - will change the inequality sign 5s<=20 s<=4

OpenStudy (anonymous):

-3\[\le\]2x+4/4

OpenStudy (ash2326):

-3<=2x+4/4 -3<=2x+1 -4<=2x -2<=x or x>=-2

OpenStudy (anonymous):

-9\[\ge\]-8(1+6v)-1

OpenStudy (anonymous):

show work plz